the pOH of a solution is 10.40. Calculate the hydrogen ion concentration in the solution. a) 4.0 x 10^-11 M b)3.6 M c)4.0 x 10^-10 d)2.5 x 10 ^-4 e)1.8 x 10^-4
Answer #1
OH concentration is 10^-10.4 = 3.98 x 10^-11
H+ conc is then 10^-14/3.98 x 10^-11 = 2.51 x 10^-4 which is choice d
Hope this helped
Answer #2

d)2.5 x 10 ^-4
Lets start by getting pH. Since pH + pOH = 14, then

pH = 14 - pOH
= 14 - 10.40
= 3.6

The formula for pH is pH = -log([H+]). So, we have

3.6 = -log([H+])
-3.6 = log([H+])
[H+] = 10^(-3.6) or about 2.5 x 10^-4 M.